Rocky Road & Company produces custom aquariums. Each aquarium passes through four departments. These four departments apply overhead based on separate rates. These rates, along with the actual consumption of cost drivers within each department, are shown below.

Department A: $100 per direct labor hour; 1,000 actual direct labor hours consumed
Department B: $75 per machine hour; 1,500 actual machine hours consumed
Department C: $0.75 per direct labor dollar; $100,000 of actual direct labor cost
Department D: $125 per maintenance work order; 500 actual maintenance work orders

The four departments also incurred the following actual overhead costs.

Department A: $110,000 actual overhead
Department B: $104,000 actual overhead
Department C: $95,000 actual overhead
Department D: $45,000 actual overhead

The firm has a policy that if overhead (that is, the total overhead for the four departments) is overapplied or underapplied by more than $5,000, the firm closes the overhead control account ratably to the four departments' WIP accounts, to the finished goods account, and to the COGS account. The firm prorates this entry using the number of units represented by each account.

Here are the units represented by each account this period.

WIP-Department A: 30 units
WIP-Department B: 10 units
WIP-Department C: 20 units
WIP-Department D: 40 units
Finished Goods: 400 units
COGS: 500 units

If overhead is overapplied or underapplied by less than $5,000, the overhead control account is just closed out to COGS.

What is the correct journal entry to close the overhead control account?

a.

Debit: Overhead Control $4,000

Credit: COGS $4,000

b.

Debit: COGS $4,000

Credit: Overhead Control $4,000

c.

Debit: COGS $52,000

Debit: FG $41,600

Debit: WIP-Department D $4,160

Debit: WIP-Department C $2,080

Debit: WIP-Department B $1,040

Debit: Dr. WIP-Department A $3,120

Credit: Overhead Control $104,000

d.

Debit: Overhead Control $104,000

Credit: COGS $52,000

Credit: FG $41,600

Credit: WIP-Department D $4,160

Credit: WIP-Department C $2,080

Credit: WIP-Department B $1,040

Credit: WIP-Department A $3,120
Answer #1

Companies generally record overhead on estimate basis and at the end of accounting period update it to actual level. Over applied overhead means overhead recorded in books of accounts are more than actual overhead. so we have to decrease it to actual level. Similarly under applied overhead means overhead recorded in books of accounts are less than actual overhead. so we have to increase overhead in books to actual level.

1.Department A

  Overhead recorded in books = Direct labor hours * overhead rate

= 1000 labor hours * 100$ per labor hour

= 100000$

Actual overhead = 110000

Since actual overhead is more than overhead recorded we can say that overhead in department A is under applied by 10000$ (Actual - recorded) (110000 - 100000)

2. Department B

  Overhead recorded in books = Direct machine hours * overhead rate

= 1500 machine hours * 75$ per machine hour

= 112500$

Actual overhead = 104000

Since actual overhead is less than overhead recorded we can say that overhead in department A is over applied by 8500$ (recorded - actual) (112500 - 104000)

3. Department C

  Overhead recorded in books = Direct labor cost * overhead rate

= 100000$ labor cost * 0.75$ per labor $

= 75000$

Actual overhead = 95000$

Since actual overhead is more than overhead recorded we can say that overhead in department A is under applied by 20000$ (Actual - recorded) (95000 -75000)

4. Department D

  Overhead recorded in books = no of Maintenance work order * overhead rate

= 500 work order * 125$ per machine hour

= 62500$

Actual overhead = 45000

Since actual overhead is less than overhead recorded we can say that overhead in department A is over applied by 17500$ (recorded - actual) (62500 - 45000)

Total over applied = Department B + Department D = 8500 + 17500 = 26000$ over applied

Total under applied = Department A + Department C = 10000 + 20000 = 30000$ under applied

Hence total overhead under applied in all four department = 30000 - 26000 = 4000

Since it is less than 5000 , so according to question overhead control account is just closed out to Cost of goods sold (COGS).

4000 under applied means we have not recorded 4000$ overhead in books till now so now we will record this and accordingly COGS will increase. So it will be a debit to COGS.

Entry will be

COGS 4000$

Overhead control 4000$

Right answer = option b